Spectrum:

Help, x+x+2+x+4+x+6=-28. State the property that justifies your first step and why you chose it

Spectrum:

Solve the linear equation.

ILikeToEatChildren:

first of all, when a variable just by itself means one, such as X=1X. so the question is saying, 1x+1x+2+1x+4+1X=28

Alexis1415:

Step 1: Add the numbers, x+x+2+x+4+x+6+=-28 x+x+12+x+x=-28

ILikeToEatChildren:

@iliketoeatchildren wrote:
first of all, when a variable just by itself means one, such as X=1X. so the question is saying, 1x+1x+2+1x+4+1X=28
Then you would add all of the numbers that have the same variable, add all numbers that have a X in the number.

Alexis1415:

Step 2: Combine like terms, 𝑥+ x+12+x+x= -28 4x+12= -28

Alexis1415:

Step 3: Subtract 12 from both sides of the equation, 4x+12= -28 4x+12-12= -28-12

Spectrum:

-40